OpenStudy (anonymous):

5. What are the real or imaginary solutions of the polynomial equation? x3 – 8 =0 (1 point)

OpenStudy (anonymous):

I have absolutely no idea how to solve this!

OpenStudy (anonymous):

\[(x-2)(x^2+2x+4)=0\] so one solution is obvious, namely \(x=2\) the other two via quadratic formula

OpenStudy (anonymous):

oh the first step was to factor \(x^3-8\) it is the difference of two cubes

OpenStudy (anonymous):

\(a^3-b^3=(a-b)(a^2+ab+b^2)\)

OpenStudy (anonymous):

1 + isquare root 3, and 1 – isquare root 3 2, –1 + isquare root 3, and –1 – isquare root 3 2, 1 + 2isquare root 3, and 1 – 2i square root 3 2, 2 +2isquare root 3, and 2 – 2isquare root 3 These are my choices. Which one is it?

OpenStudy (anonymous):

do you know the quadratic formula?

OpenStudy (anonymous):

Sorta.

OpenStudy (anonymous):

that is what you need

OpenStudy (anonymous):

\[x^2+2x+4=0\] use \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=1,b=2, c=4\]

OpenStudy (anonymous):

actually in this case it is easier to complete the square

OpenStudy (anonymous):

\[x^2+2x+4=0\] \[x^2+2x=-4\] \[(x+1)^2=-4+1=-3\] \[x+1=\pm\sqrt{-3}\] \[x=-1\pm\sqrt{-3}\]

OpenStudy (anonymous):

Thanks so much! :)

OpenStudy (anonymous):

yw